# SusBench: An Online Benchmark for Evaluating Dark Pattern Susceptibility of Computer-Use Agents

### v1

@@ 1. Introduction
As large language model (LLM)-powered computer-use agents (CUAs) increasingly show potential to assist human users in tasks like shopping, information retrieval, and schedule to-do lists (OpenAI, 2025; Wang et al., 2025a), users expect agents to interact with user interfaces in ways that align with their intent. This means navigating a complex web environment replete with ambiguities that can sometimes render it difficult for users themselves to act in accordance with their intentions. Users are particularly likely to fail to act in their own best interests when the digital environment is engineered to subtly manipulate or deceive them into taking actions that benefit the provider of a web service. In the field of human-computer interaction, such user interface (UI) designs are referred to as “dark patterns” (Gray et al., 2018) or “deceptive design patterns” (Chordia et al., 2023; Deceptive Patterns, 2025), and they present a unique challenge for CUAs: how to act in accordance with user intention across long web trajectories, even as interfaces along that trajectory seek to undermine this goal.
Previous work in fields ranging from human-computer interaction (Mildner et al., 2023) to psychology (Waldman, 2020) to marketing (Runge et al., 2023) has demonstrated that humans are vulnerable to a wide array of dark patterns. Even though such patterns can seem obvious when they are explicitly identified, they are often easy to miss in real-world contexts. For example, an interface may preselect a protection plan or subscription service by default, making it easy for users to opt into such services (and potentially pay for them or agree to exchange additional data with a provider) unless they consciously opt out (Mathur et al., 2021a). If CUAs are vulnerable to dark patterns, their ability to perform online tasks on the user’s behalf may be compromised. Conversely, capable CUAs could be better at identifying and avoiding dark patterns than human users, making them potentially excellent delegates for carrying out tasks on users’ behalf. Thus, in this study, we sought to understand: RQ1) whether CUAs are vulnerable to user interface dark patterns and, RQ2) how, if at all, they differ from human users.
@@ 2.2. Dark Patterns and AI
Additionally, past work has found that LLMs can generate interface code that contains unsolicited dark patterns (Krauß et al., 2025; Chen et al., 2025a). For example, Krauß et al. (2025) found that generative coding models (GPT-4) consistently create websites that include dark patterns without being prompted by the developer to do so, and with no warning provided to the developer. In this work, we utilize this property of LLMs to generate dark pattern injections, which allow us to construct evaluation tasks of dark pattern UI at scale.
@@ 2.3. AI Agents, Computer-Use Agents, and Benchmarks
AI agents are LLM-based systems that can plan out and take their own actions (Yao et al., 2023). They are trained to perform tasks like using digital tools (Schick et al., 2023), interacting with websites and APIs (Nakano et al., 2021), challenging their own reasoning (Zhou et al., 2025b), and collaborating with other agents (Zhou et al., 2025a), soliciting direction from a user as needed. In the present work, we focus on computer-use agents (CUAs), which operate a computer as a human might, for example, by interacting with file stores, operating systems, and, importantly for this study, websites and user interfaces (Agashe et al., 2024; Wang et al., 2025a). CUAs are typically equipped with both vision capabilities and reasoning capabilities that allow them to observe the visual and textual elements of a website, reflect on the actions available to them, and carry out actions according to the intentions of a user (OpenAI, 2025).
@@
Recent work identified several limitations of CUAs relevant to the present work. Ma et al. (2024) found that multimodal LLMs are susceptible to distraction in information-rich environments, resulting in misalignment with user intentions. Chen et al. (2025b) found that LLM-based GUI agents are susceptible to fine-print injections, especially to contextually embedded threats. Moreover, many scholars have noted that CUAs and other agents may deceptively “fake” alignment with the user, appearing to carry out the user’s intention while in fact pursuing other ends (Greenblatt et al., 2024; Ji et al., 2025). Our work is complementary to this research, in that we study the deceptive engineering of the environment navigated by an agent, which can result in failure to accord with a user’s intent.
@@
To better measure CUAs’ performance on navigating through interfaces and performing actions on behalf of humans, much prior work has attempted to benchmark the performance of these agents across long-horizon tasks. These benchmarks can be categorized into two types: offline and online benchmarks. Offline benchmarks, such as Mind2Web and WebLINX (Deng et al., 2023; Lù et al., 2024), operate by caching portions of real-world websites, enabling rapid iteration and reproducibility during agent development, but this setting inherently lacks completeness due to the dynamic and interactive nature of live web environments (Xue et al., 2025a). Online benchmarks, on the other hand, offer realistic and interactive environments that best approximate how CUAs are actually used. Many online benchmarks have been developed to evaluate CUAs. He et al. (2024) introduced WebVoyager to evaluate CUAs on 15 live, real-world websites across 812 tasks. Zhou et al. (2023) created WebArena, which contains sandbox environments for evaluating the capabilities of autonomous AI agents. Xue et al. (2025a) introduced the Online-Mind2Web benchmark, built on top of the original Mind2Web (Deng et al., 2023), to measure CUAs’ capabilities with 135 real websites, using an automated LLM-based judge. To test how agents perform in malicious environments, Boisvert et al. (2025) introduced DoomArena, a framework which allows injecting malicious content into the user-agent-environment loop to test agents’ vulnerability to security threats. We build on this large body of prior work by focusing on evaluating agents’ susceptibility to more subtle deceptive designs in naturalistic and online settings, using a code injection method.
@@ 5.2.2. Dark Pattern Avoidance Rate
Human |
| Overall | 67.5 | 27.7 | 68.3 | 33.1 | 63.8 | 35.4 | 66.0 | 32.0 | 63.7 | 29.9 | 62.4 | 34.1 | ||||||||||||
| Confirm Shaming | 85.3 | 6.1 | 93.9 | 4.2 | 91.2 | 4.9 | 87.5 | 5.8 | 82.9 | 6.4 | 97.1 | 2.8 | ||||||||||||
| Disguised Ad | 60.0 | 8.9 | 69.0 | 8.6 | 45.2 | 8.9 | 46.7 | 9.1 | 83.9 | 6.6 | 83.9 | 6.6 | ||||||||||||
| Fake Social Proof | 90.9 | 5.0 | 93.3 | 4.6 | 93.9 | 4.2 | 93.9 | 4.2 | 92.6 | 5.0 | 93.3 | 4.6 | ||||||||||||
| False Hierarchy | 100.0 | 0.0 | 100.0 | 0.0 | 100.0 | 0.0 | 100.0 | 0.0 | 95.2 | 4.6 | 100.0 | 0.0 | ||||||||||||
| Forced Action | 91.7 | 4.6 | 91.4 | 4.7 | 91.2 | 4.9 | 91.7 | 4.6 | 84.8 | 6.2 | 82.9 | 6.4 | ||||||||||||
| Hidden Information | 17.6 | 6.5 | 8.6 | 4.7 | 8.3 | 4.6 | 8.6 | 4.7 | 14.7 | 6.1 | 9.4 | 5.2 | ||||||||||||
| Pop-Up Ad | 86.2 | 6.4 | 100.0 | 0.0 | 96.6 | 3.4 | 93.1 | 4.7 | 65.5 | 8.8 | 65.5 | 8.8 | ||||||||||||
| Preselection | 44.1 | 8.5 | 36.7 | 8.8 | 13.3 | 6.2 | 33.3 | 9.1 | 24.1 | 7.9 | 18.8 | 6.9 | ||||||||||||
| Trick Wording | 46.7 | 7.4 | 42.9 | 7.6 | 51.2 | 7.6 | 52.3 | 7.5 | 44.1 | 8.5 | 34.1 | 7.1 | ||||||||||||
@@ 6.2. Potential Use of Computer-Use Agents as Simulations of Human Users in Dark Pattern Evaluation
This finding suggests that CUAs might be helpful for supporting automatic evaluation of user interface designs, particularly when assessing the impact of manipulative designs. For researchers, such automatic evaluation can be valuable given the challenges of conducting similar human studies. For example, testing dark pattern susceptibility with real humans may be unethical if the pattern exploits people’s psychological and emotional vulnerabilities (Mathur et al., 2021b). In such cases, CUAs could serve as the first line of evaluation and allow for more cautious, subsequent human experiments. For user experience designers, synthetic users simulated through CUAs can be used to support early-stage testing of user interfaces. They can potentially help flag manipulative designs before their deployment. Similarly, CUAs’ comparable sensitivity to manipulative design cues also suggests that they can serve as useful tools in detecting and auditing deceptive design practices in online services. For example, a CUA can be used to systematically traverse through a website’s user flow. Using the trajectories it produces, auditors can identify potential manipulative designs and understand how it may influence real-world users. By leveraging CUAs for large-scale, reproducible testing, designers, researchers and regulators could more efficiently identify risky UI configurations, benchmark compliance with design standards, and iteratively refine interfaces to promote more transparent and trustworthy user experiences.
However, we caution that CUAs’ similarity to human susceptibility does not guarantee behavioral fidelity across diverse contexts. Agents lack emotions, fatigue, and variable attention that shape real human interactions. Using LLM-based agents as proxies for human participants may also come with its own risks. While past studies have found promising results of using agents to simulate human behavior (Park et al., 2023, 2022, 2024), researchers have argued that replacing human participants can harmfully misportray identity groups (Wang et al., 2025) and can raise ethical and epistemological concerns (Kapania et al., 2025). Others also showed more nuanced evidence, indicating that LLMs can replicate certain results of psychological experiments but not others (Cui et al., 2024). While CUAs may serve as a meaningful proxy for human users, data derived from CUAs should be taken critically and be validated against real human data.
@@ 6.3. Implications for Regulation and Policymaking
Looking ahead, as the web becomes increasingly mediated by autonomous CUAs, regulatory frameworks will need to evolve beyond human-centered consumer protection. Agents will interact with interfaces at scale, potentially amplifying or mitigating the harms of dark patterns depending on how they are trained and aligned. Yet, this shift also blurs the boundaries of responsibility. When an agent interacts with a deceptive interface and takes an unintended action, such as making a purchase or accepting a subscription, it becomes unclear where accountability lies. The responsibility may rest with the website designer who deployed the manipulative pattern, or the developer of the agent that failed to detect it. There are already many efforts underway to define the future of AI governance, such as the proposed “Safe and Secure Innovation for Frontier Artificial Intelligence Models Act” in California (Legislature, 2024), the European Union’s AI Act (Parliament, 2025), and Colorado’s AI Act (Rodriguez et al., 2024). Our results suggest that such legislation should incorporate guidance for a world where agents move about the web—shopping, posting comments, curating information, and more—on humans’ behalf. Future regulation may need to consider not only the design of user interfaces but also the behavioral standards, transparency mechanisms, and auditability of autonomous agents, ensuring that both sides of the interaction—interfaces and agents—operate under shared expectations of fairness, predictability, and accountability.
@@ 6.4. Limitations
This study has a few limitations. First, demographics are limited to well-educated young people with significant shopping experience. Their susceptibility might differ form other populations, which we did not evaluate in this present work. Second, since all participants finished the study in a controlled lab setting, it is unknown how much their behavior resembles real-world shopping behavior. Third, our evaluation only includes a binary outcome, and future work may explore more nuanced measurement of dark pattern avoidance, such as by considering user’s explicit and implicit intent.

### v2

@@ 1. Introduction
As large language model (LLM)-powered computer-use agents (CUAs) increasingly show potential to assist human users in tasks like shopping, information retrieval, and scheduling to-do lists (OpenAI, 2025; Wang et al., 2025c), users expect agents to interact with user interfaces in ways that align with their intent. This means navigating a complex web environment replete with ambiguities that can sometimes render it difficult for users themselves to act in accordance with their intentions. Users are particularly likely to fail to act in their own best interests when the digital environment is engineered to subtly manipulate or deceive them into taking actions that benefit the provider of a web service. In the field of human-computer interaction, such user interface (UI) designs are referred to as “dark patterns” (Gray et al., 2018) or “deceptive design patterns” (Chordia et al., 2023; Deceptive Patterns, 2025), and they present a unique challenge for CUAs: how to act in accordance with user intention across long web trajectories, even as interfaces along that trajectory seek to undermine this goal.
Previous work in fields ranging from human-computer interaction (Mildner et al., 2023) to psychology (Waldman, 2020) to marketing (Runge et al., 2023) has demonstrated that humans are vulnerable to a wide array of dark patterns. Even though such patterns can seem obvious when they are explicitly identified, they are often easy to miss in real-world contexts. For example, an interface may preselect a protection plan or subscription service by default, making it easy for users to opt into such services (and potentially pay for them or agree to exchange additional data with a provider) unless they consciously opt out (Mathur et al., 2021b). If CUAs are vulnerable to dark patterns, their ability to perform online tasks on the user’s behalf may be compromised. Conversely, capable CUAs could be better at identifying and avoiding dark patterns than human users, making them potentially excellent delegates for carrying out tasks on users’ behalf. Thus, in this study, we sought to understand: RQ1) whether CUAs are vulnerable to user interface dark patterns and, RQ2) how, if at all, they differ from human users.
@@ 2.2. Dark Patterns and AI
Additionally, past work has found that LLMs can generate interface code that contains unsolicited dark patterns (Krauß et al., 2025a; Chen et al., 2025c). For example, Krauß et al. (2025b) found that generative coding models (GPT-4) consistently create websites that include dark patterns without being prompted by the developer to do so, and with no warning provided to the developer. In this work, we utilize this property of LLMs to generate dark pattern injections, which allow us to construct evaluation tasks of dark pattern UI at scale.
@@ 2.3. AI Agents, Computer-Use Agents, and Benchmarks
AI agents are LLM-based systems that can plan out and take their own actions (Yao et al., 2023). They are trained to perform tasks like using digital tools (Schick et al., 2023), interacting with websites and APIs (Nakano et al., 2022), challenging their own reasoning (Zhou et al., 2025b), and collaborating with other agents (Zhou et al., 2025a), soliciting direction from a user as needed. In the present work, we focus on computer-use agents (CUAs), which operate a computer as a human might, for example, by interacting with file stores, operating systems, and, importantly for this study, websites and user interfaces (Agashe et al., 2024; Wang et al., 2025c). CUAs are typically equipped with both vision capabilities and reasoning capabilities that allow them to observe the visual and textual elements of a website, reflect on the actions available to them, and carry out actions according to the intentions of a user (OpenAI, 2025).
@@
Recent work identified several limitations of CUAs relevant to the present work. Ma et al. (2025) found that multimodal LLMs are susceptible to distraction in information-rich environments, resulting in misalignment with user intentions. Chen et al. (2025a) found that LLM-based GUI agents are susceptible to fine-print injections, especially to contextually embedded threats. Moreover, many scholars have noted that CUAs and other agents may deceptively “fake” alignment with the user, appearing to carry out the user’s intention while in fact pursuing other ends (Greenblatt et al., 2024; Ji et al., 2025). Our work is complementary to this research, in that we study the deceptive engineering of the environment navigated by an agent, which can result in failure to accord with a user’s intent.
@@
To better measure CUAs’ performance on navigating through interfaces and performing actions on behalf of humans, much prior work has attempted to benchmark the performance of these agents across long-horizon tasks. These benchmarks can be categorized into two types: offline and online benchmarks. Offline benchmarks, such as Mind2Web and WebLINX (Deng et al., 2023; lù2024weblinxrealworldwebsitenavigation), operate by caching portions of real-world websites, enabling rapid iteration and reproducibility during agent development, but this setting inherently lacks completeness due to the dynamic and interactive nature of live web environments (Xue et al., 2025a). Online benchmarks, on the other hand, offer realistic and interactive environments that best approximate how CUAs are actually used. Many online benchmarks have been developed to evaluate CUAs. He et al. (2024) introduced WebVoyager to evaluate CUAs on 15 live, real-world websites across 812 tasks. Zhou et al. (2024) created WebArena, which contains sandbox environments for evaluating the capabilities of autonomous AI agents. Xue et al. (2025a) introduced the Online-Mind2Web benchmark, built on top of the original Mind2Web (Deng et al., 2023), to measure CUAs’ capabilities with 135 real websites, using an automated LLM-based judge. To test how agents perform in malicious environments, Boisvert et al. (2025) introduced DoomArena, a framework which allows injecting malicious content into the user-agent-environment loop to test agents’ vulnerability to security threats. We build on this large body of prior work by focusing on evaluating agents’ susceptibility to more subtle deceptive designs in naturalistic and online settings, using a code injection method.
@@ 3.4. Protocol for Creating Dark Pattern Injections
SusBench supports injecting both static dark patterns (i.e., dark patterns with fixed UI elements) and dynamic ones (i.e., dark patterns that dynamically change based on webpage content such as item information on a shopping website). The difference lies in whether the generated injection constructs only hard-coded elements or dynamic elements selected from the webpage (e.g., through CSS selectors). While dynamic dark patterns may be more reflective of real-world dark patterns in certain cases, incorporating overly dynamic dark patterns could introduce confounding variables to the evaluation. For example, changes in text or placement of dark pattern can potentially impact agent behavior, adding noise to evaluation results. Thus, in our existing evaluation set, we ensure that all evaluation tasks are largely consistent (especially the design, layout, and writing that are core to the dark pattern), and only include small tweaks to the text to if they are needed to better match the webpage content. For example, the injected dark pattern text may contain a substring extracted from the item information, and prices associated with injected additional purchases can change based on the original price of the item.
@@ 3.6. Benchmark Implementation
This figure presents a system diagram of SusBench, illustrating how dark pattern injections are deployed and evaluated during task execution. The diagram is laid out horizontally, with five main components connected by directional arrows that indicate data flow and control. Each component is represented as a rounded rectangle with an icon and label. Four labeled interaction points—(A) through (D)—correspond to stages described in the caption.
[…]
• Injection Function Store (top left, green): A database-like component that stores pre-generated dark pattern injection functions, each associated with an injection ID.
• Browser Extension (top center-right, green): A browser-side component responsible for injecting dark patterns, matching webpages, and evaluating outcomes.
• Controller (bottom left, blue): A central orchestration component that selects injections, assigns tasks, and collects evaluation results.
• Human / Agent (bottom center, gray): Represents either a human participant or an autonomous agent performing tasks in the browser.
• Playwright Browser (right, red): A controlled browser environment where webpages are loaded and interacted with.
[…]
An arrow from the Controller points upward to the Injection Function Store, labeled “Injection ID,” indicating that the Controller randomly selects an injection ID. The corresponding injection functions are retrieved and sent to the Browser Extension. After a dark pattern is injected and encountered, the Browser Extension applies an evaluation function to determine the outcome (e.g., whether the dark pattern was avoided) and sends the evaluation result back to the Controller.
[…]
An arrow labeled “Task” flows from the Controller to the Human / Agent. This indicates that the Controller presents a task description—such as a shopping or navigation goal—that the Human or Agent is instructed to complete in the browser.
[…]
The Browser Extension connects to the Playwright Browser via arrows labeled “Page Match & Inject” and “Input Events.” When the Playwright Browser loads a webpage, the Browser Extension checks whether the page matches an injection target and, if so, injects the dark pattern using the selected injection functions. Input events generated during interaction (such as clicks or form submissions) are sent back from Playwright to the Browser Extension.
[…]
Bidirectional arrows connect the Human / Agent and the Playwright Browser, labeled “Observe” and “Act.” The Human or Agent receives webpage observations rendered by Playwright and responds by acting on the browser through inputs such as mouse clicks or keyboard actions.
[…]
@@ 4.1.1. Agent and Model Selection
All the models we used except GPT-4o are reasoning models with “thinking mode” turned on for agentic tasks. All the agents we used generate reasoning data in the final trajectories.
@@ 5.2.2. Dark Pattern Avoidance Rate
| Dark Pattern Avoidance Rate (%) | Human | BU (Gpt-5) | BU (Gemini-2.5-Pro) | BU (Claude-Sonnet-4) | CUA (Gpt-4o) | CU (Claude-Sonnet-4) | ||||||
| Overall | 67.5 | 27.7 | 68.3 | 33.1 | 63.8 | 35.4 | 66.0 | 32.0 | 63.7 | 29.9 | 62.4 | 34.1 |
| Confirm Shaming | 85.3 | 6.1 | 93.9 | 4.2 | 91.2 | 4.9 | 87.5 | 5.8 | 82.9 | 6.4 | 97.1 | 2.8 |
| Disguised Ad | 60.0 | 8.9 | 69.0 | 8.6 | 45.2 | 8.9 | 46.7 | 9.1 | 83.9 | 6.6 | 83.9 | 6.6 |
| Fake Social Proof | 90.9 | 5.0 | 93.3 | 4.6 | 93.9 | 4.2 | 93.9 | 4.2 | 92.6 | 5.0 | 93.3 | 4.6 |
| False Hierarchy | 100.0 | 0.0 | 100.0 | 0.0 | 100.0 | 0.0 | 100.0 | 0.0 | 95.2 | 4.6 | 100.0 | 0.0 |
| Forced Action | 91.7 | 4.6 | 91.4 | 4.7 | 91.2 | 4.9 | 91.7 | 4.6 | 84.8 | 6.2 | 82.9 | 6.4 |
| Hidden Information | 17.6 | 6.5 | 8.6 | 4.7 | 8.3 | 4.6 | 8.6 | 4.7 | 14.7 | 6.1 | 9.4 | 5.2 |
| Pop-Up Ad | 86.2 | 6.4 | 100.0 | 0.0 | 96.6 | 3.4 | 93.1 | 4.7 | 65.5 | 8.8 | 65.5 | 8.8 |
| Preselection | 44.1 | 8.5 | 36.7 | 8.8 | 13.3 | 6.2 | 33.3 | 9.1 | 24.1 | 7.9 | 18.8 | 6.9 |
| Trick Wording | 46.7 | 7.4 | 42.9 | 7.6 | 51.2 | 7.6 | 52.3 | 7.5 | 44.1 | 8.5 | 34.1 | 7.1 |
@@ 6. Discussion
Below, we discuss the implications of our findings for computer-use agent design, the use of agents as proxies for human users, and the regulation of dark patterns in online interfaces.
@@ 6.2. Potential Use of Computer-Use Agents as Simulations of Human Users in Dark Pattern Evaluation
This finding suggests that CUAs might be helpful for supporting automatic evaluation of user interface designs, particularly when assessing the impact of manipulative designs. For researchers, such automatic evaluation can be valuable given the challenges of conducting similar human studies. For example, testing dark pattern susceptibility with real humans may be unethical if the pattern exploits people’s psychological and emotional vulnerabilities (Mathur et al., 2021a). In such cases, CUAs could serve as the first line of evaluation and allow for more cautious, subsequent human experiments. For user experience designers, synthetic users simulated through CUAs can be used to support early-stage testing of user interfaces. They can potentially help flag manipulative designs before their deployment. Similarly, CUAs’ comparable sensitivity to manipulative design cues also suggests that they can serve as useful tools in detecting and auditing deceptive design practices in online services. For example, a CUA can be used to systematically traverse through a website’s user flow. Using the trajectories it produces, auditors can identify potential manipulative designs and understand how it may influence real-world users. By leveraging CUAs for large-scale, reproducible testing, designers, researchers and regulators could more efficiently identify risky UI configurations, benchmark compliance with design standards, and iteratively refine interfaces to promote more transparent and trustworthy user experiences.
However, we caution that CUAs’ similarity to human susceptibility does not guarantee behavioral fidelity across diverse contexts. Agents lack emotions, fatigue, and variable attention that shape real human interactions. Using LLM-based agents as proxies for human participants may also come with its own risks. While past studies have found promising results of using agents to simulate human behavior (Park et al., 2023, 2022, 2024), researchers have argued that replacing human participants can harmfully misportray identity groups (Wang et al., 2025a) and can raise ethical and epistemological concerns (Kapania et al., 2025). Others also showed more nuanced evidence, indicating that LLMs can replicate certain results of psychological experiments but not others (Cui et al., 2024). While CUAs may serve as a meaningful proxy for human users, data derived from CUAs should be taken critically and be validated against real human data.
@@ 6.3. Implications for Regulation and Policymaking
Looking ahead, as the web becomes increasingly mediated by autonomous CUAs, regulatory frameworks will need to evolve beyond human-centered consumer protection. Agents will interact with interfaces at scale, potentially amplifying or mitigating the harms of dark patterns depending on how they are trained and aligned. Yet, this shift also blurs the boundaries of responsibility. When an agent interacts with a deceptive interface and takes an unintended action, such as making a purchase or accepting a subscription, it becomes unclear where accountability lies. The responsibility may rest with the website designer who deployed the manipulative pattern, or the developer of the agent that failed to detect it. There are already many efforts underway to define the future of AI governance, such as the proposed “Safe and Secure Innovation for Frontier Artificial Intelligence Models Act” in California (Legislature, 2024), the European Union’s AI Act (Parliament, 2025), and Colorado’s AI Act (Rodriguez et al., 2024). Our results suggest that such legislation should incorporate guidance for a world where agents move about the web—shopping, posting comments, curating information, and more—on humans’ behalf. Future regulation may need to consider not only the design of user interfaces but also the behavioral standards, transparency mechanisms, and auditability of autonomous agents, ensuring that both sides of the interaction (interfaces and agents) operate under shared expectations of fairness, predictability, and accountability.
@@ 6.4. Limitations
This study has a few limitations. First, demographics are limited to well-educated young people with significant shopping experience. Their susceptibility might differ form other populations, which we did not evaluate in this present work. Thus, the claims made about human users’ susceptibility to dark patterns in this paper should be interpreted with caution. More future studies are needed to definitively understand the difference between human and agent users’ dark pattern susceptibility. Second, since all participants finished the study in a controlled lab setting, it is unknown how much their behavior resembles real-world shopping behavior. Third, our evaluation only includes a binary outcome, which can be a simplification of the effects of dark patterns. This binary approach may also be limited for evaluating the kinds of dark patterns that exploit user’s attention and thus do not require a specific action from the user, such as infinite scrolling and autoplay (Chen et al., 2025b; Schaffner et al., 2025). Future work may explore more nuanced measurement of dark pattern avoidance, such as by considering user’s explicit and implicit intent, and incorporating the effects of attention-prolonging dark patterns. Finally, a risk that comes with the dark pattern generation pipeline and the publicly available benchmark in this paper is that these tools could be exploited by website owners and third parties to insert dark patterns into existing pipelines. To mitigate such possibilities, our tool requires the user to explicitly opt into installing the browser extension on their browser (if it were to be used standalone without the isolated Playwright environment). The tool itself also does not support inserting code into an existing codebase, but rather creates another layer of injections into the browser runtime environment, thus preventing easy integration of dark patterns into existing websites or services.